# A Comprehensive Survey of Deep Learning for Multivariate Time Series Forecasting: A Channel Strategy Perspective

### v2

@@ 4 Comparison within the Taxonomy
In this section, we compare the strengths and limitations of CI, CD, and CP across multiple dimensions—see Table 2.
Efficiency: Efficiency measures the amount of resources consumed by a model during its operation, such as time and memory. CI is the most efficient strategy as it processes each channel independently without modeling inter-channel relationships. This results in the lowest computational complexity and excellent scalability for large datasets. CD is the least efficient strategy as it requires modeling all inter-channel dependencies. The computational complexity increases sharply with the number of channels, making it less scalable. The CP strategy strikes a balance by dynamically capturing interactions among channels while allowing each channel to focus only on the ones most relevant to itself. Thanks to its dynamical mechanism, which limits the modeling scope, CP remains more efficient than CD.
Robustness: Robustness refers to the ability of a model to maintain stability and effectiveness against noise, data variations, or interference. CI has a certain degree of robustness to noise as it processes each channel independently, avoiding interference among channels. CD has higher robustness for strongly correlated channels but is highly sensitive to noise. Spurious correlations can significantly degrade its performance, making it the least robust strategy. CP exhibits the high robustness by capturing flexible and dynamic relationships, making it highly effective in handling noise and variations in data distribution.
Generalizability: Generalizability refers to the ability of a model to perform well on unseen data or different datasets by leveraging patterns and relationships beyond the training data. CI has the weakest generalizability as it cannot leverage channel correlations, which is a critical drawback for multivariate time series tasks. CD demonstrates strong generalizability when channel correlations are consistent and significant. However, its performance deteriorates when relationships are weak or vary significantly across datasets. CP exhibits the strongest generalizability by handling overlapping and dynamic channel correlations. It adapts well to various datasets, particularly in complex real-world scenarios.
Capacity: Capacity refers to the ability of a model to capture and represent complex relationships and dependencies within the data. CI has the lowest capacity as it completely ignores inter-channel relationships and can only model the dynamics of individual channels. CD has the highest capacity as it simultaneously models all inter-channel relationships, allowing it to capture complex global dependencies. The CP strategy offers a balanced approach with moderate capacity. It models interactions selectively, allowing each channel to focus only on the channels most relevant to it.
Ease of Implementation: Ease of Implementation refers to how straightforward or complex it is to put a model into practice, considering the required components and design. CI is the easiest strategy to implement due to its simple structure, as it does not require modeling inter-channel relationships. CD is more complex to implement because it requires designing modules to capture inter-channel dependencies. CP is the most challenging strategy to implement as it requires a dynamic and flexible mechanism to model inter-channel dependencies. Typically, this involves the use of attention mechanisms or graph-based methods, which add to the difficulty of implementation.
@@ 5.3 Multi-modality for Channel Correlations
Multiple modalities can be introduced to more comprehensively model the correlation among channels. Compared to a single time series modality, multimodal data can provide richer information sources, such as text, images, or other sensor data, which can compensate for potential gaps in time series data. By extracting features from multimodal data, the unique characteristics of channels across different modalities can be captured. Subsequently, cross-modal relational modeling mechanisms, such as cross-modal attention mechanisms or GNN, can be employed to uncover the dynamic dependencies among channels. Additionally, to further enhance the modeling of channel correlation, an adaptive fusion mechanism can be designed to dynamically adjust interaction weights based on the correlations among different modalities.
@@ 5.4 Channel Strategy of Foundation Models
Multivariate time series foundation models follow two main approaches: LLM-based models and time series pre-trained models. LLM-based models, lacking a channel dimension in language modality, typically adopt a CI strategy Jin et al. (2024); Chang et al. (2023). Due to the high heterogeneity in the number of channels in time series data, most time series pre-trained models, such as Timer Liu et al. (2024c) and Chronos Ansari et al. (2024), use a CI strategy to ensure robust predictions while avoiding complex channel correlation. In contrast, models like MOIRIA Woo et al. (2024) and UniTS Gao et al. (2024) incorporate channel correlation during pretraining. MOIRIA flattens all channels, using positional embeddings to distinguish them, capturing both temporal and channel relationships with self-attention, while UniTS directly captures channel correlations via self-attention in the channel dimension. However, time series pre-trained models only consider CD strategy to capture the channel correlations, without fully considering the intricate and diverse channel correlations in different pretain datasets, where CP strategy may achieve better performance. There also lacks works in LLM-based models to consider the channel correlations combined with the multimodal data. Existing approaches remain relatively basic, leaving significant room for improving channel strategies in foundation models.

### v3

@@ 4 Comparison within the Taxonomy
This section compares the advantages and limitations of CI, CD, and CP from multiple dimensions. We select three representative models for each category. To evaluate these three-channel strategies, we conduct two experiments.
Performance across Correlation: Figure 6 presents the MSE using a lookback window of 512 and a prediction horizon of 96. The experimental data is drawn from the time series forecasting benchmark, TFB Qiu et al. (2024), from which nine datasets are selected to represent varying correlation strengths. For visual clarity, model families are color-coded: CI (blue), CD (orange), and CP (green). In weakly correlated scenarios (left), CI achieves superior predictive performance by naturally disregarding inter-channel noise, whereas CD exhibits degraded performance due to its tendency to overfit spurious correlations. However, as correlations strengthen (right), CI encounters a capacity bottleneck, causing its performance to lag behind that of CD and CP strategies. In these regimes, both CD and CP excel, with the CP strategy demonstrating the strongest generalization. By flexibly modeling inter-channel dependencies, CP effectively captures complex interactions while filtering noise, yielding superior performance compared to pure CD strategies.
Efficiency and Computational Complexity: Figure 7 details the inference time and model parameter counts using a batch size of 1, a lookback window of 512, and a prediction horizon of 96. We specifically select the ETTm2 and PEMS-BAY datasets as representatives of scenarios with minimum and maximum inter-channel correlations, respectively. The CI strategy (purple boxed region) adopts a channel-independent paradigm, decoupling multivariate time series into univariate processes and enabling weight sharing across all channels. This design effectively eliminates the computational redundancy associated with inter-channel interactions, thereby minimizing parameter counts and significantly reducing inference latency to achieve superior computational efficiency. In stark contrast, CD and CP strategies (red boxed region), driven by the need to explicitly model complex inter-channel dependencies, inevitably incorporate additional architectural components (such as attention mechanisms or dynamic graph structures). This architectural complexity results in a substantial computational burden, manifested as significantly higher parameter counts and inference costs compared to the CI strategy.
@@ 5.3 Multi-modality for Channel Correlations
Multiple modalities can be introduced to more comprehensively model the correlation among channels. Compared to a single time series modality, multimodal data can provide richer information sources, such as text, images, or other sensor data, which can compensate for potential gaps in time series data. By extracting features from multimodal data, the unique characteristics of channels across different modalities can be captured. Subsequently, cross-modal relational modeling mechanisms, such as cross-modal attention mechanisms or GNNs, can be employed to uncover the dynamic dependencies among channels. Additionally, to further enhance the modeling of channel correlation, an adaptive fusion mechanism can be designed to dynamically adjust interaction weights based on the correlations among different modalities.
@@ 5.4 Channel Strategy of Foundation Models
Multivariate time series foundation models follow two main approaches: LLM-based models and time series pre-trained models. LLM-based models, lacking a channel dimension in language modality, typically adopt a CI strategy Jin et al. (2024). Due to the high heterogeneity in the number of channels in time series data, most time series pre-trained models, such as Timer Liu et al. (2024c) , use a CI strategy to ensure robust predictions while avoiding complex channel correlation. In contrast, models like MOIRIA Woo et al. (2024) and UniTS Gao et al. (2024) incorporate channel correlation during pretraining. MOIRIA flattens all channels, using positional embeddings to distinguish them, capturing both temporal and channel relationships with self-attention, while UniTS directly captures channel correlations via self-attention in the channel dimension. However, time series pre-trained models only consider the CD strategy to capture the channel correlations, without fully considering the intricate and diverse channel correlations in different pre-training datasets, where the CP strategy may achieve better performance. There is also lacks works in LLM-based models to consider the channel correlations combined with the multimodal data. Existing approaches remain relatively basic, leaving significant room for improving channel strategies in foundation models.